I'm not sure on how to solve this equation:
f(x)=4x+7 and g(x)= 3x-5 find (gxf)(-4)

Answer by stanbon(75887)

f(x)=4x+7 and g(x)= 3x-5 find (gxf)(-4)
------
(g*f)(x) = (4x+7)(3x-5) = 12x^2 + x -35
-----
(g*f)(-4) = 12(-4)^2 + (-4) - 35 = 12*16 -39 = 153
------------
Cheers,
Stan H.
-----------